OpenStudy (anonymous):

The limit represents f '(c) for a function f(x) and a number c. Find f(x) and c. lim [7 (sqrt x) -21] / (x-9) x-> 9

OpenStudy (irishboy123):

\[\large \lim_{{x \to 9}} \frac{7 \sqrt x -21}{x-9}\] hit it with l''Hospital's rule

OpenStudy (badhi):

\[\frac{7(\sqrt x - 3)}{(\sqrt x^2 -3^2)}\]

OpenStudy (irishboy123):

nicer